1.     A cob Douglas production function for a firm is given as Q=4L ½ K½. The firm has also established that wage rate and interest paid on capital are $3 and $5 respectively for a production period. The firm intents to spend $200 million for the period on production cost. Compute the levels of capital and labor that will maximize output. What is the maximum output?



1
Expert's answer
2021-03-02T17:44:35-0500
"\frac {\delta Q}{\delta L}=2(\frac{K}{L})^{0.5}"

"\frac {\delta Q}{\delta K}=2(\frac{L}{K})^{0.5}"

"3L+5K=200"

"L=0.6K"

"K=29"

"L=17"